OpenStudy (anonymous):

The difference of two number is 2. What is the smallest possible value of their product?

OpenStudy (ankitshaw):

if u take the two numbers as 1 and -1 then the product is -1

OpenStudy (ankitshaw):

but it will be positive and u need the smallest product

OpenStudy (acxbox22):

sorry forgot that simple rule for a sec :D

OpenStudy (anonymous):

x - y = 2 Solving for y we get y = x - 2 The product = xy = x ( x - 2) = x^2 - 2x To maximize that take the derivative y' = 2x - 2 Set = 0 and sove for x 2x - 2 = 0 x = 1 y = -1 The smallest product is 1 * -1 which equals -1

OpenStudy (anonymous):

thnx.. :D